# Deepfake Detection that Generalizes Across Benchmarks

### v1

@@ Abstract
The generalization of deepfake detectors to unseen manipulation techniques remains a challenge for practical deployment. Although many approaches adapt foundation models by introducing significant architectural complexity, this work demonstrates that robust generalization is achievable through a parameter-efficient adaptation of a pre-trained CLIP vision encoder. The proposed method, LNCLIP-DF, fine-tunes only the Layer Normalization parameters (0.03% of the total) and enhances generalization by enforcing a hyperspherical feature manifold using L2 normalization and latent space augmentations.
We conducted an extensive evaluation on 13 benchmark datasets spanning from 2019 to 2025. The proposed method achieves state-of-the-art performance, outperforming more complex, recent approaches in average cross-dataset AUROC. Our analysis yields two primary findings for the field: 1) training on paired real-fake data from the same source video is essential for mitigating shortcut learning and improving generalization, and 2) detection difficulty on academic datasets has not strictly increased over time, with models trained on older, diverse datasets showing strong generalization capabilities.
This work delivers a computationally efficient and reproducible method, proving that state-of-the-art generalization is attainable by making targeted, minimal changes to a pre-trained CLIP model. The code will be made publicly available upon acceptance.
@@ 1 Introduction
The generalization gap is the primary issue that we address in this work. Assuming the hypothesis that adapted large-scale, pre-trained vision model can serve as a general foundation for deepfake detection [35], we select the Contrastive Language-Image Pre-training (CLIP) [38] model as a feature extractor, which is known for its generalizable visual representations.
The proposed method consists of a CLIP’s vision encoder, whose outputs are L2 normalized. We then fine-tune only the parameters of the Layer Normalization blocks [37] while keeping the rest of CLIP frozen. In addition to that, we propose using metric learning in this L2 space, together with latent augmentation, to improve generalization.
We benchmarked the generalization capabilities of the proposed model on 13 deepfake video datasets released between 2019 and 2025, listed in  Tab. 1. To our knowledge, this represents the broadest evaluation in the deepfake literature. We show that the proposed model outperforms the most recent state-of-the-art methods on the majority of all available benchmarks.
@@
  • A novel deepfake detection method called LNCLIP-DF. The method achieves the best average cross-dataset AUROC compared to recently released models.
@@ 3 Method
The proposed method uses the pre-trained CLIP ViT-L/14 image encoder, to which we add three components: L2 normalization of the classification token, slerp augmentation [42], and a linear classifier. We optimize only 0.03% of network parameters using a weighted combination of cross-entropy, uniformity, and alignment losses in an end-to-end fashion. We call this approach LNCLIP-DF, and its complete overview is shown in Fig. 1.
Figure 1: The architecture of LNCLIP-DF. The gray rectangle represents the original CLIP ViT image encoder, and green represents the added components. Only rectangles with a fire icon represent layers with trainable parameters. The slerp block is used only during training.
@@ 3.1 Model
Feature extractor.
We use the CLIP ViT-L/14 image encoder, which was trained by OpenAI [38]. The model output is $16\times 16+1$ tokens of dimension $1024$, where $16\times 16$ represents the number of visual patches and $1$ refers to the classification token. We use only the classification token and discard all patch tokens.
L2 normalization.
The classification token is L2 normalized. Restricting features to live within a hypersphere has been observed to help achieve better results in face-related tasks, such as face recognition [40, 11]. Discarding the information on the magnitudes of the feature vectors encourages the model to create clusters within hyperspherical caps. From the perspective of a linear classifier, classification is akin to separating a hyperspherical cap from the rest of a hypersphere [48].
Optimized parameters.
The L2-normalized classification token is used as the input to the linear binary classifier, which gives logits for fake and real classes. In addition to the weights $W\in\mathbb{R}^{1024\times 2}$ and biases $b\in\mathbb{R}^{2}$ of the classification layer, we optimize the parameters of all Layer Normalization blocks [37, 16, 47]. This roughly constitutes 0.03% of the total model parameters; the remaining parameters of the CLIP are frozen.
Slerp.
Similarly to [53], we augment the data in the latent feature space. As we constrained the features to hyperspherical space, we developed another approach to augment the latent space. We are using spherical linear interpolation (slerp) [42] to interpolate between features of the same class within the same batch. For each L2-normalized feature $\mathbf{z}_{i}$ in a batch, we randomly sample a feature $\mathbf{z}_{j}$ from the same batch that has the same class $y_{i}=y_{j}$. We also sample an interpolation parameter $t\sim\mathcal{U}(0,1)$ from the uniform distribution. The slerp is
$\text{slerp}(\mathbf{z}_{i},\mathbf{z}_{j};t)=\frac{\sin((1-t)\theta)}{\sin\theta}\mathbf{z}_{i}+\frac{\sin(t\theta)}{\sin\theta}\mathbf{z}_{j}\;,$ (1)
where $\theta=\arccos(\mathbf{z}_{i}^{\intercal}\mathbf{z}_{j})$, with $\|\mathbf{z}_{i}\|=\|\mathbf{z}_{j}\|=1$. This augmentation extends the batch. For the fake class, it helps fill the latent space between feature clusters formed by samples from generators present in the training data. We hypothesize that this can potentially prepare the model to detect samples from new generators whose forgery fingerprints may resemble some combination of features from already seen generators.
Loss function.
The network is trained using a weighted combination of cross-entropy, uniformity, and alignment losses [48]
@@
Training algorithm.
To optimize the model parameters, we use Adam optimizer [27] with $\beta_{1}=0.9$ and $\beta_{2}=0.999$ without weight decay. The weight precision is set to bfloat16. The learning rate is scheduled using a cosine cyclic rule [43]. Each cycle starts with a linear warm-up for one epoch, increasing from $10^{-5}$ to $3\times 10^{-4}$, and then decays using cosine scheduling over nine epochs to $10^{-5}$. For most runs, we observed that the training did not improve after two cycles (or 20 epochs). The batch size is set to 128 samples. The loss coefficients of Eq. 2 are set to $\alpha=0.1$ and $\beta=0.5$. The slerp augmentation extends the batch to 1024 samples.
Video classification.
The proposed model has a single frame as an input. To get video-level probabilities, we take 32 frames sampled from every video, calculate softmax probabilities for every frame independently, and average them.
@@ 3.2 Data
Data preprocessing.
We follow a standard practice of dataset preprocessing proposed in DeepfakeBench [52], which consists of the following steps:
@@
    Crop and resize the face to $256\times 256$.
@@
Training data.
Following the standard protocol, we train the model on the FaceForensics++ dataset [39] (FF++), which consists of $3600$ videos, of which $720$ are real videos, and the rest, $4\times 720$, come from four different deepfake forgery methods. After preprocessing, we obtain approximately 115115k frames. The dataset was released in three different compression levels. Following all of the preceding work, we use the c23 compression version of the dataset. We augment the training set using image augmentations such as random horizontal flipping, random affine transformations, Gaussian blurring, color jitter, and JPEG compression, all implemented in the torchvision package. In ablations, we also experimented with changing the training set to FFIW [60] and DSv2 [3].
Training samples pairing.
We train the model on the dataset which consists of real-fake pairs, where each fake video is generated from the real counterpart. We hypothesize and empirically confirm in Sec. 4.4 that the dataset constructed in this way forces the network to learn the subtle, low-level artifacts of the manipulation process, rather than exploiting superficial high-level differences between unrelated videos, thus preventing shortcut learning.
Validation data.
We noticed that the FF++ validation set is very similar to the training set. That makes the comparison of different training experiments on such a validation ineffective. Each experiment achieves the best validation AUROC quickly, shows almost no signs of overfitting, and does not provide any estimate of cross-dataset generalization. The FF++ validation set does not provide crucial information on how the methods will perform when trained for an extended period. Therefore, we create a custom validation set, which consists of data from validation or training splits of other datasets, namely CDFv2 [30], DSv1 and DSv2 [3], and FFIW [60]. We ensure that test samples from these datasets are not present in the custom validation set.
@@ 4 Experiments
Table 1: Test datasets. The number of real and fake videos in the datasets. Negative values: the number of videos missing in our experiments, compared to the original datasets, due to face detector failures. Values in rows with datasets marked * represent the number of videos randomly subsampled from the original dataset.
Year Dataset Real Fake
2019 FF++ [39] 140 560
2019 DF [39] 140 140
2019 F2F [39] 140 140
2019 FS [39] 140 140
2019 NT [39] 140 140
2019 DFD [14] 363 3068 -2
2019 UADFV [56] 49 49
2019 DFDC [13] 2500 -185 2500 -111
2020 CDFv2 [30] 178 340
2021 FFIW [60] 1738 -3 1738 -3
2021 KoDF [28] * 403 1106
2021 FakeAVCeleb [25] 500 20566 -22
2022 DFDM [24] 590 -2 1720 -2
2024 PolyGlotFake [22] 762 13605
2024 IDForge [50] * 18834 $23$
2024 DeepSpeak v1.1 [3] 1416 1497
2025 DeepSpeak v2.0 [3] 1863 1416
We demonstrate the generalizability of LNCLIP-DF by presenting a series of experiments. First, we introduce the diverse range of benchmark datasets and the evaluation metric. Next, we conduct a comprehensive cross-dataset evaluation, where we train the model on the FF++ dataset and compare its performance against numerous state-of-the-art (SOTA) methods on other benchmarks. Following this, a detailed ablation study is presented to systematically analyze the contribution of each component of the proposed method. We then investigated two hypotheses: one demonstrating the critical importance of training on paired real-fake data for robust generalization and another exploring the evolution of deepfake detection difficulty over the years, which shows that training on older, diverse datasets can be more beneficial than using solely recent datasets.
@@ 4.1 Test benchmarks
For reporting the performance of the model in cross-dataset settings, we are using Celeb-DF-v2 (CDFv2) [30], DeepFake Detection Challenge (DFDC) [13], Google’s DFD dataset [14], Face Forensics in the Wild (FFIW) [60], DeepSpeak v1.1 (DSv1) and DeepSpeak v2.0 (DSv2) [3], Korean DeepFake Detection Dataset (KoDF) [28], FakeAVCeleb (FAVC) [25], DeepFakes from Different Models (DFDM) [24], PolyGlotFake (PGF) [22], and IDForge (IDF) [50]. Dataset statistics are shown in  Tab. 1.
We are using the video-level area under the ROC curve (AUROC) as the main comparison and optimization metric.
@@ 4.2 Cross-dataset evaluation
Table 2: Video-level AUROC in cross-dataset testing of models trained on the FF++ dataset. Results of other methods are taken from their original papers. Values with superscript citations are taken from the papers referenced in superscripts.
Model Year Publication CDFv2 DFD DFDC FFIW
LipForensics [17] 2021 CVPR 82.4 73.5
@@
RealForensics [18] 2022 CVPR 86.9 75.9
SBI [41] 2022 CVPR 93.2 82.7 72.4 84.8
AUNet [58] 2023 CVPR 92.8 99.2 73.8 81.5
StyleDFD [8] 2024 CVPR 89.0 96.1
@@
LAA-Net [33] 2024 CVPR 95.4 98.4 86.9
AltFreezing [49] 2024 CVPR 89.5 98.5
NACO [57] 2024 ECCV 89.5 76.7
RAE [46] 2024 ECCV 95.5 99.0 80.2
TALL++ [51] 2024 IJCV 92.0 78.5
OPR [7] 2024 NeurIPS 92.5 77.0
UDD [15] 2025 AAAI 93.1 95.5 81.2
P&P [55] 2025 CVPR 94.7 96.5 84.3 92.1
DFD-FCG [19] 2025 CVPR 95.0 81.8
ForAda [9] 2025 CVPR 95.7 97.2 87.2
Effort [54] 2025 ICML 95.6 96.5 84.3 92.1
LNCLIP-DF (proposed) 2025 96.5 97.3 87.0 93.5
We compared the proposed model in a cross-dataset fashion following the standard protocol, i.e., by training on FF++ and testing on the rest of the datasets as in previous work. Table 2 compares the video-level AUROC of the proposed model with state-of-the-art (SOTA) approaches. AUROC values for the competing methods were taken directly from the original papers. The model achieves SOTA results on two datasets: CDFv2 and FFIW, and gets high results on the DFDC dataset.
The proposed method achieves these results simply by averaging individual frame softmax probabilities, without complex post-processing or architectural modifications as described in Sec. 3.1. This contrasts with many contemporary methods that rely on complex multimodal analyses, such as modeling temporal inconsistencies between video frames or detecting audio-video discrepancies [17, 32, 18, 59, 19, 55]. The performance of the proposed method suggests that the features of the CLIP visual encoder, when properly tuned, are highly effective for this detection task.
We also evaluated the generalization of the proposed method against current SOTA models across a wide range of benchmark datasets. For this experiment, we selected the most recent open-source SOTA models that have made their code and pre-trained weights publicly available, namely ForAda [9] and Effort [54]. To ensure that each model was used as intended by its original authors, we utilized their officially released weights and integrated their data preprocessing pipelines which were used during model training. After reproducing the results on the test datasets that were used in these papers, we obtained the same or very similar results, presented in Tab. 4.
@@
Table 3: Comparison of methods based on video-level test AUROC in %, recalculated by us. In-dataset: the test and training sets are i.i.d. Cross-dataset: test set includes data from generators not represented in the training set. The highest score in each column is in bold.
In-dataset Cross-dataset
@@
ForAda [9] 96.8 99.7 97.0 98.6 91.9 99.4 97.2 87.3 82.0 95.7 90.6 88.2 93.1 97.1 86.6 90.8 81.8 72.8 91.5
Effort [54] 93.9 99.4 93.2 98.4 84.6 97.4 95.2 85.4 91.2 93.2 92.5 88.1 92.4 98.2 84.9 96.0 82.1 64.4 90.6
LNCLIP-DF 98.0 99.6 98.2 98.6 95.7 99.8 97.3 87.0 88.2 96.5 93.5 85.9 96.3 98.8 92.9 97.3 92.6 80.0 94.2
Table 4: Reimplementation of ForAda [9] and Effort [54] deepfake detectors. Reported / Reproduced Video-Level AUROC (%).

### v2

@@ Abstract
The generalization of deepfake detectors to unseen manipulation techniques remains a challenge for practical deployment. Although many approaches adapt foundation models by introducing significant architectural complexity, this work demonstrates that robust generalization is achievable through a parameter-efficient adaptation of one of the foundational pre-trained vision encoders. The proposed method, GenD, fine-tunes only the Layer Normalization parameters (0.03% of the total) and enhances generalization by enforcing a hyperspherical feature manifold using L2 normalization and metric learning on it.
We conducted an extensive evaluation on 14 benchmark datasets spanning from 2019 to 2025. The proposed method achieves state-of-the-art performance, outperforming more complex, recent approaches in average cross-dataset AUROC. Our analysis yields two primary findings for the field: 1) training on paired real-fake data from the same source video is essential for mitigating shortcut learning and improving generalization, and 2) detection difficulty on academic datasets has not strictly increased over time, with models trained on older, diverse datasets showing strong generalization capabilities.
This work delivers a computationally efficient and reproducible method, proving that state-of-the-art generalization is attainable by making targeted, minimal changes to a pre-trained foundational image encoder model. The code will be made publicly available upon acceptance.
@@ 1 Introduction
The generalization gap is the primary issue that we address in this work. Assuming the hypothesis that adapted large-scale, pre-trained foundational vision encoder can serve as a general foundation for deepfake detection [40], we build the proposed method in three variants, using Contrastive Language-Image Pre-training (CLIP) [43], Perception Encoder (PE) [8] and DINO [46] models as feature extractors, which are known for their generalizable visual representations.
The proposed method consists of a vision encoder, whose outputs are L2-normalized. We then fine-tune only the parameters of the Layer Normalization blocks [42] while keeping the rest frozen. Additionally, we propose using metric learning in this L2 space to enhance generalization.
We benchmarked the generalization capabilities of the proposed model on 14 deepfake video datasets released between 2019 and 2025, listed in  Tab. 1. To our knowledge, this represents the broadest evaluation in the deepfake literature. We show that the proposed model outperforms the most recent state-of-the-art methods on the majority of all available benchmarks.
@@
  • A novel deepfake detection method called GenD. The method achieves the best average cross-dataset AUROC compared to recently released models.
@@ 2 Related Work
As a complimentary approach to the passive detection methods of AI manipulations, there exist active defense approaches that protect real images by incorporating invisible watermarks [65, 3, 6, 62, 2, 61].
@@ 3 Method
The proposed method comes in three variants depending on the pre-trained image encoder to which we add two components: L2 normalization of the classification token and a linear classifier. We optimize only 0.03% of network parameters using a weighted combination of cross-entropy, uniformity, and alignment losses in an end-to-end fashion. We call this approach GenD, and its complete overview is shown in Fig. 1.
Figure 1: The architecture of GenD (CLIP). The gray rectangle represents the original CLIP ViT image encoder, and green represents the added components. Only rectangles with a fire icon represent layers with trainable parameters.
@@ 3.1 Model
Feature extractor. We use three different pre-trained image encoders: 1. CLIP ViT-L/14, trained by OpenAI [43]; 2. $\text{PE}_{\text{core}}\text{L}$, trained by Meta AI [8]; 3. DINOv3 ViT-L/16 [46], trained by Meta AI. After the encoder processes the image, we take only the $1024$-dimensional classification token and discard the rest.
Optimized parameters. The L2-normalized classification token is used as the input to the linear binary classifier, which gives logits for fake and real classes. In addition to the weights $W\in\mathbb{R}^{1024\times 2}$ and biases $b\in\mathbb{R}^{2}$ of the classification layer, we optimize the parameters of all Layer Normalization blocks [42, 19, 51]. This roughly constitutes 0.03% of the total model parameters; the remaining parameters of the encoder are frozen.
Loss function. The network is trained using a weighted combination of cross-entropy, uniformity, and alignment losses [52]
@@
Training algorithm. To optimize the model parameters, we use Adam optimizer [30] with $\beta_{1}=0.9$ and $\beta_{2}=0.999$ without weight decay. The weight precision is set to bfloat16. The learning rate is scheduled using a cosine cyclic rule [47]. Each cycle starts with a linear warm-up for one epoch, increasing from $10^{-5}$ to $3\times 10^{-4}$, and then decays using cosine scheduling over nine epochs to $10^{-5}$. For most runs, we observed that the training did not improve after two cycles (or 20 epochs). The batch size is set to 128 samples. For CLIP ViT-L/14, the loss coefficients of Eq. 1 are set to $\alpha=0.1$ and $\beta=0.5$, found on the validation set.
Video classification. GenD has a single frame as an input. To get video-level probabilities, we uniformly sample 32 frames from a video, calculate softmax probabilities for every frame independently, and average them.
Speed and memory. Every encoder of GenD is of the size of a ViT-L model, consisting of roughly 304M parameters. Our non-optimized for speed and memory implementation in PyTorch can process 120 frames per second on A100 with a batch size of 1, requiring approximately 2GB of VRAM.
@@ 3.2 Data
Data preprocessing. We follow a standard practice of dataset preprocessing proposed in DeepfakeBench [56], which consists of the following steps:
@@
    Crop and resize the face to $224\times 224$.
@@
Training data. Following the standard protocol, we train the model on the FaceForensics++ dataset [44] (FF++), which consists of $3600$ videos, of which $720$ are real videos, and the rest, $4\times 720$, come from four different deepfake forgery methods. After preprocessing, we obtain approximately 115115k frames. The dataset was released in three different compression levels. Following all of the preceding work, we use the c23 compression version of the dataset. We augment the training set using image augmentations such as random horizontal flipping, random affine transformations, Gaussian blurring, color jitter, and JPEG compression, all implemented in the torchvision package. In ablations, we also experimented with changing the training set to FFIW [67] and DSv2 [5].
Training samples pairing. We train the model on a dataset consisting of real-fake pairs, where each fake video is generated from its real counterpart. We hypothesize and empirically confirm in Sec. 4.4 that the dataset constructed in this way forces the network to learn the subtle, low-level artifacts of the manipulation process, rather than exploiting superficial, high-level differences between unrelated videos, thus preventing shortcut learning.
Validation data. We noticed that the FF++ validation set is very similar to the training set. That makes the comparison of different training experiments on such a validation ineffective. Each experiment achieves the best validation AUROC quickly, shows almost no signs of overfitting, and does not provide any estimate of cross-dataset generalization. The FF++ validation set does not provide crucial information on how the methods will perform when trained for an extended period. Therefore, we create a custom validation set, which consists of data from validation or training splits of other datasets, namely CDFv3 [35], DSv1 and DSv2 [5], and FFIW [67]. We ensure that test samples from these datasets are not present in the custom validation set.
@@ 4 Experiments
Table 1: Test datasets. The number of real and fake videos in the datasets. Negative values: the number of videos missing in our experiments, compared to the original datasets, due to face detector failures. Values in rows with datasets marked * represent the number of videos randomly subsampled from the original dataset. Gen. means the number of generators used in the dataset.
Year Dataset Gen. Real Fake
2019 FF++ [44] 4 140 560
2019 DF [44] 1 140 140
2019 F2F [44] 1 140 140
2019 FS [44] 1 140 140
2019 NT [44] 1 140 140
2019 DFD [17] 5 363 3068 -2
2019 UADFV [60] 1 49 49
2019 DFDC [16] 8 2500 -185 2500 -111
2020 FSh [32] 1 140 140
2020 CDFv2 [34] 1 178 340
2021 FFIW [67] 3 1738 -3 1738 -3
2021 KoDF [31] * 6 403 1106
2021 FAVC [28] 4 500 20566 -22
2022 DFDM [27] 5 590 -2 1720 -2
2024 PGF [25] 10 762 13605
2024 IDF [54] * 9 18834 $23$
2024 DSv1.1 [5] 5 1416 1497
2025 DSv2 [5] 6 1863 1416
2025 CDFv3 [35] 22 178 2540 -1
We demonstrate the generalizability of GenD by presenting a series of experiments. First, we introduce the diverse range of benchmark datasets and the evaluation metric. Next, we conduct a comprehensive cross-dataset evaluation, where we train the model on the FF++ dataset and compare its performance against numerous state-of-the-art (SOTA) methods on other benchmarks. Following this, a detailed ablation study is presented to systematically analyze the contribution of each component of the proposed method. We then investigated two hypotheses: one demonstrating the critical importance of training on paired real-fake data for robust generalization and another exploring the evolution of deepfake detection difficulty over the years, which shows that training on older, diverse datasets can be more beneficial than using solely recent datasets. We finish the experiments with robustness evaluation against common image degradation techniques.
We ensure the statistical significance of the results by repeating each experiment 5 times with different training seeds. Every reported AUROC for GenD except for the Tab. 2 is averaged over 5 seeds; standard deviations can be found in the supplementary text.
@@ 4.1 Test benchmarks
For reporting the performance of the model in cross-dataset settings, we are using Celeb-DF-v2 (CDFv2) [34], Celeb-DF++ (CDFv3) [35], DeepFake Detection Challenge (DFDC) [16], Google’s DFD dataset [17], Face Forensics in the Wild (FFIW) [67], DeepSpeak v1.1 (DSv1.1) and DeepSpeak v2.0 (DSv2) [5], Korean DeepFake Detection Dataset (KoDF) [31], FakeAVCeleb (FAVC) [28], DeepFakes from Different Models (DFDM) [27], PolyGlotFake (PGF) [25], and IDForge (IDF) [54]. The statistics for the test part of the datasets are shown in  Tab. 1.
We are using the video-level area under the ROC curve (AUROC) as the main comparison and optimization metric. See the supplementary material for more evaluation metrics such as average precision (AP) or equal error rate (EER).
@@ 4.2 Cross-dataset evaluation
Table 2: Video-level AUROC (%) in cross-dataset testing of models trained on the FF++ dataset. Results of other methods are taken from their original papers. Values with superscript citations are taken from the papers referenced in superscripts. Video input means that the model takes a sequence of frames and models temporal correlations between them.
Model Year Publication Input Backbone CDFv2 DFD DFDC FFIW
LipForensics [20] 2021 CVPR Video ResNet-18 82.4 73.5
@@
RealForensics [21] 2022 CVPR Video Modified CSN 86.9 75.9
SBI [45] 2022 CVPR Frame EFNB4 93.2 82.7 72.4 84.8
AUNet [64] 2023 CVPR Video Xception+ART 92.8 99.2 73.8 81.5
StyleDFD [12] 2024 CVPR Video 3D ResNet-50 89.0 96.1
@@
LAA-Net [38] 2024 CVPR Frame EFNB4 95.4 98.4 86.9
AltFreezing [53] 2024 CVPR Video 3D ResNet-50 89.5 98.5
NACO [63] 2024 ECCV Video ViT-B/16 89.5 76.7
RAE [50] 2024 ECCV Frame ViT-B/16 95.5 99.0 80.2
TALL++ [55] 2024 IJCV Video Swin-B 92.0 78.5
ProDet [11] 2024 NeurIPS Frame EFNB4 92.5 77.0
UDD [18] 2025 AAAI Frame CLIP ViT-B/16 93.1 95.5 81.2
P&P [59] 2025 CVPR Video CLIP ViT-L/14 94.7 96.5 84.3 92.1
DFD-FCG [22] 2025 CVPR Video CLIP ViT-L/14 95.0 81.8
ForAda [13] 2025 CVPR Frame CLIP ViT-L/14 95.7 97.2 87.2
Effort [58] 2025 ICML Frame CLIP ViT-L/14 95.6 96.5 84.3 92.1
GenD (CLIP) 2025 Frame CLIP ViT-L/14 96.0 97.0 87.1 92.8
GenD (PE) 2025 Frame $\text{PE}_{\text{core}}\text{L}$ 95.8 96.5 81.6 93.3
GenD (DINO) 2025 Frame DINOv3 ViT-L/16 92.2 96.6 84.7 94.5
We compared the proposed model in a cross-dataset fashion following the standard protocol, that is, by training on FF++ and testing on the rest of the datasets as in previous work. Tab. 2 compares the video-level AUROC of the proposed model with SOTA approaches. AUROC values for the competing methods were taken directly from the original papers. The model achieves SOTA results on two datasets: CDFv2 and FFIW, and gets high results on the DFDC dataset.
The proposed method achieves these results simply by averaging individual frame softmax probabilities, without complex post-processing or architectural modifications as described in Sec. 3.1. This contrasts with many contemporary methods that rely on complex multimodal analyses, such as modeling temporal inconsistencies between video frames or detecting audio-video discrepancies [20, 37, 21, 66, 22, 59]. The performance of the proposed method suggests that the features of the visual encoder, when properly tuned, are highly effective for this detection task.
We also evaluated the generalization of the proposed method against current SOTA models across a wide range of benchmark datasets. For this experiment, we selected the most recent open-source SOTA models that have made their code and pre-trained weights publicly available, namely ForAda [13] and Effort [58]. To ensure that each model was used as intended by its original authors, we utilized their officially released weights and integrated their data preprocessing pipelines which were used during model training. After reproducing the results on the test datasets that were used in these papers, we obtained the same or very similar results, presented in Tab. 5. The in-distribution results on FF++ are shown in Tab. 3.
@@
Table 3: Video-level test AUROC (%) on in-domain FF++ dataset calculated by us. GenD results are the means over five seeds.
Method DF F2F FS NT Mean
ForAda [13] 99.7 97.0 98.6 91.9 96.8
Effort [58] 99.4 93.2 98.4 84.6 93.9
GenD (CLIP) 99.5 98.1 98.7 95.5 98.0
GenD (PE) 99.7 99.3 99.1 97.5 98.9
GenD (DINO) 99.6 99.4 99.4 96.7 98.8
Table 4: Cross-dataset video-level AUROC (%) for reproduced methods. The highest score in each column is in bold. Results for GenD are the averages over five training seeds; standard deviations can be found in the supplementary.
@@
ForAda [13] 99.4 97.2 87.3 82.0 95.7 90.6 88.2 93.1 97.1 86.6 90.8 81.8 72.8 75.6 88.4
Effort [58] 97.4 95.2 85.4 91.2 93.2 92.5 88.1 92.4 98.2 84.9 96.0 82.1 64.4 78.7 88.5
GenD (CLIP) 99.2 96.4 86.4 86.6 94.6 91.5 84.9 96.0 99.6 89.6 97.8 90.1 77.7 85.9 91.2
GenD (PE) 97.7 96.8 82.2 87.6 95.0 93.7 85.1 97.3 98.3 92.3 97.9 87.8 78.6 89.5 91.4
GenD (DINO) 98.6 96.2 85.6 88.8 92.5 92.9 89.7 98.4 99.8 92.4 98.2 86.9 79.4 83.5 91.6
Table 5: Comparison of Reported / Reproduced video-level AUROC (%) for reimplemented deepfake detectors.